OpenStudy (faman39):

The volume of a cube is 144 cubic meters. What is the volume of the square based pyramid with the same side length and height? 16 m3 48 m3 432 m3 1, 296 m3

OpenStudy (faman39):

@baldymcgee6 can you able to help me ?

OpenStudy (baldymcgee6):

What is the formula for the volume of a cube?

OpenStudy (faman39):

ah, 1/2 ?

OpenStudy (baldymcgee6):

The volume of a cube is: V = s^3

OpenStudy (faman39):

last one, right?

OpenStudy (faman39):

No, i think i m wrong

OpenStudy (baldymcgee6):

you are wrong.. :) v = s^3 144 = s^3 s = cbrt(144) V_pyramid = (area of base)(height)*(1/3)

OpenStudy (faman39):

So you think it is third one ?

OpenStudy (baldymcgee6):

nope..

OpenStudy (faman39):

So... i dont really get this one

OpenStudy (baldymcgee6):

that's okay.. i will draw it up for you

OpenStudy (ghazi):

 does that help fatima? @Faman39

OpenStudy (faman39):

No :(

OpenStudy (ghazi):

oops wait